1.) Find the area between the curve f(x) = \{ \begin{matrix}2 & x < 0 \\ −3 & x > 0 \end{matrix} , and the x-axis, and between x = −4 and x = 5.

2a.) ∫_{−4}^{5} |f(x)|dx =

2b.) ∫_{−4}^{5} f(x)dx = 3.) The speed of a runner decreased steadily after crossing the finish line. Her speed at 2 second intervals is given in the table. Find lower and upper estimates for the distance that she traveled during these 6 seconds.

\begin{matrix}t(seconds) & 0 & 2 & 4 & 6 \\ v(feet/sec) & 40 & 20 & 5 & 0 \end{matrix}

Lower estimate: , Upper estimate: